If the ratio of the ages of Kissi and Esinam is 3:5 and that of Esinam and Lariba is 3:5 and the sum of the ages of all 3 is 147 years, what is the age difference between oldest the youngest?
1
Expert's answer
2020-06-10T18:54:01-0400

Given: K/E= 3/5, E/L=3/5, K+L+E=147

Solution:

Rewrite K =3E/5 ----------------Equation 1

Rewrite L=5E/3 ------------------Equation 2

K+L+E=147 ------------------------Equation 3


Substitute Equation 1 and 2 in equation 3


3E/5 +5E/3 + E= 147


(9E+25E+15E)/15= 147


49E= 2205


E= 2205/49=45


Age of Esinam=45 years

Age of Kissi=3(45)/5=27 years

Age of Lariba=5(45)/3=75 years


Hence difference between oldest and youngest = 75-27= 48 years.